OpenStudy (anonymous):

solve this quadratic equation by completing the square. x2+2x=20 x=1 plus or minus square root of 24 x=-1 plus or minus square root of 24 x=-1 plus or minus square root of 21 x=1 plus or minus square root of 21

OpenStudy (calculusxy):

1. \[x^2 + 2x = 20 \rightarrow x^2 + 2x - 20 = 20 - 20\] 2. \[ x^2 + 2x - 20 =0 \rightarrow x^2 + 2x - 20 + 20 = 0 + 20\] 3. \[x^2 + 2x = 20 \rightarrow x^2 + 2x + (2 \div 1) = 20 + (2 \div 1)\] 4. \[x^2 + 2x + 1 = 20 + 1 \rightarrow \sqrt{(x + 1)^2} = \sqrt{21} \rightarrow (x+1) \pm \sqrt{21}\]

OpenStudy (anonymous):

which "completes" the square? Hint: only one later takes square root of a perfect square

OpenStudy (anonymous):

so is it x= -1 plus or minus square root of 21?

OpenStudy (anonymous):

should be x+1 = \(\pm \sqrt{21}\) solve for x